Commissioning Notes for 20080423 UT

Summary:

LBTtools/IRTC
  • center script allows interactive positioning of source on detector
  • takepic, takeseq, and takeEFpair keep track of lastimg previmg, lastseq, prevseq (todo)

Found approximate offset for AGW on-axis (or off)

Found AGW patrol field is SOUTH of IRTC when at PA=0

First test of LBTI SW interface to IIF/TCS

Running Notes/Log (UTtimes):

Open at sunset. New SW build had problems starting up, fix in progress.

03:15 Bringing up TCS GUIs... OK, but HBS down.

03:32 pointto M5_0444 - failed (HBS died right about the time the preset was sent)

03:32 Starting up AGW software. Check4nport OK, oacserver seems OK, AGW1 started (waiting for UMAC to warm past 10.00C)

03:49 Sending pointto M5_0444 again. Note PA=0 near SE starts at low rotator angles and will track to limit at the meridian crossing.

pointto M5_0441 in NE, PA=0 should be OK. pointto WT10_363 need fainter star to take movie of oscillations. Oscillations are not clear now.

Testing GCS. Stages moved after a minor bug fix. Guiding initiated, not clear it is seeing a real source.

04:58 Polynomial problem, followed by an HBS failure.

Two other HBS failures in rapid succession.

Determined the approximate offsets in their stage coordinate system to move the detector on-axis (ideally, PCS should give SFP coordinates of the on-axis source near 0,0)
  • Xstage(mm) = Xpcs(mm) + 7.5
  • Ystage(mm) = 600.5 - Ypcs(mm)

Determined that the AGW patrol field is SOUTH of the IRTC at PA=0

Vidya did some initial testing of the LBTI software, will fix some bugs and retest tomorrow.

Fixed POSANGLE problem in fits headers.

center script now offsets a source to the rotator center at any PA. The normal use is:
  • pointto
  • takepic
  • center
  • use interactive mode to move source somewhere other than the rotator center.
  • Will include taking the picture as part of a future upgrade, as well as a final image as sanity check and a switch on whether the offsets are relative or absolute.

started modifying takepic, takeseq, takeEFpair to keep track of the current image (lastimg) as well as the one before that (previmg)

11:00 (the "4:00am problem"?) Rotator not receiving polynomials. Then the mount polys stopped. We could not recover from this in ~15 minutes, so we shut down a little early to take some darks in a dark dome.

11:38 taking some 245K darks (the other day's were at 270), H filter, dark dome
  • 114544 & 00005_cube (takeseq 0.10 100)
  • 114655 & 00006_cube (takeseq 0.25 100)
  • 114804 & 00007_cube (takeseq 0.50 100)
  • 114931 & 00008_cube (takeseq 1.00 100)
  • 115157 & 00009_cube (takeseq 0.01 100)
  • 115251 & 00010_cube (takeseq 0.05 100)
  • 115500 & 00011_cube (takeseq 0.001 100)

12:04 repeat darks in J filter, dark dome
  • 121027 & 00017_cube (takeseq 0.10 100)
  • 121121 & 00018_cube (takeseq 0.25 100)
  • 121214 & 00019_cube (takeseq 0.50 100)
  • 121334 & 00020_cube (takeseq 1.00 100)
  • 121547 & 00021_cube (takeseq 0.01 100)
  • 121616 & 00022_cube (takeseq 0.05 100)
  • 121718 & 00023_cube (takeseq 0.001 100)

-- DavidThompson - 23 Apr 2008
Topic revision: r3 - 23 Apr 2008, DavidThompson
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback